# Edge of Stochastic Stability: Revisiting the Edge of Stability for SGD

### v3

@@ Appendix F On Largest Eigenvalues of Sums of Matrices
F.1 Ordering the Largest Eigenvalues.
@@
F.2 Trends of $\lambda_{\max}$ given $b$
@@
As we discuss in Section 5.3, while the behavior we observe for very small batches is not surprising, for bigger batch sizes it is. In computer vision tasks where there are way more parameters than datapoints, we observe that the gap between $\lambda_{\max}^{b}$ and $\lambda_{\max}$ decreases linearly with $1/b$ when evaluating the $\lambda_{\max}^{b}$ of any fixed model with different batch sizes, see Figure 8. This $1/b$ scaling is also what we expect from our mathematical analysis in Appendix F. We are indeed able to establish that for a fixed net, the gap scales as $1/b$ for small batch sizes, when the worst case $\lambda_{\max}^{b}$ is very different from $\lambda_{\max}$, and with $1/\sqrt{b}$ for big batch sizes [^2] Although big $b$ means such that the number of directions spanned in the parameter space by the vectors $\nabla_{\theta}f(\theta,x)$ are repeated multiple times, and that may be practically unrealistic with the current sizes of networks.. Indeed:
@@
where $\sigma^{2}=\frac{1}{b}\max\{\|\mathbb{E}[\sum_{i}M_{i}^{\top}M_{i}\|,\|\mathbb{E}[\sum_{i}M_{i}M_{i}^{\top}\|\}$ is the expected second moment of the matrices and $B\geq\|M_{i}-M\|$, for all $i$, is a bound to the biggest random matrix $M_{i}$.
F.3 Random Matrix Theory for Scaling Eigenvalues.
@@
Let $n_{1},n_{2},b\in\mathbb{N}$, let $M_{1},M_{2},\ldots,M_{b}\in\mathbb{R}^{n_{1}\times n_{2}}$ be independent random matrices satisfying $\mathbb{E}[M_{i}]=M$ and $\|M_{i}-M\|\leq B$ for all $i$, let $v=\max\{\|\mathbb{E}[\sum_{i}M_{i}^{\top}M_{i}\|,\|\mathbb{E}[\sum_{i}M_{i}M_{i}^{\top}\|\}$ then for all $t>0$
@@
Appendix G Dependence of $\lambda_{\max}^{b}$-$\lambda_{\max}$ Gap on the Batch Size
@@ G.2 The trained case
Appendix H The Hessian and the Fisher Information Matrix Overlap
@@
Appendix I Exemplification Through a Simplified Model
@@
  • Smaller size: Thus setting $\lambda_{\max}$ equal to $\lambda$ means that the max between $\lambda_{1}$ and $\lambda_{2}$ is exactly $2\lambda$. Note that the fact that $a_{1}\cdot b_{1}=a_{2}\cdot b_{2}=1$ and Cauchy-Schwartz imply that $\lambda_{1},\lambda_{2}\geq 2$. Setting $\lambda_{\max}^{1}$ to $\lambda$ thus implies that the maximum between $\lambda_{1}$ and $\lambda_{2}$ is at most $22\lambda-$ 2, generally smaller.
@@
Appendix J Illustration of EoSS in Variety of Settings: Batch Sharpness
@@
We further confirm the EoSS regime in a five-layer CNN. As depicted in Figures 24 and 18, Batch Sharpness continues to plateau near the instability threshold for two distinct step sizes, while $\lambda_{\max}$ once again settles at a lower level. Notably, as we vary the batch size, the gap between Batch Sharpness and $\lambda_{\max}$ increases for smaller batches, mirroring the patterns described in Section 5.
@@
Figure 17: MLP: 2 hidden layers, hidden dimension 512; learning rate 0.01, 8k subset of CIFAR-10. Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue dots).
@@
Figure 18: CNN: 5 layers (3 convolutional, 2 fully-connected), learning rate 0.03, 8k subset of CIFAR-10. Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue dots).
@@
Figure 19: ResNet-10, learning rate 0.005, 8k subset of CIFAR-10. Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue dots).
Appendix K Illustration of EoSS in Variety of Settings: $\lambda_{\max}^{b}$
@@
To assess whether increased depth alters the phenomenon, we use a deeper MLP (MLP_L) with four hidden layers, training again on the same CIFAR-10 subset. Figures 22 and 23 show that $\lambda_{\max}^{b}$ exhibits the same EoSS behavior for two different step sizes, reinforcing that depth alone does not invalidate our findings.
@@
We further confirm the EoSS regime in a five-layer CNN. As depicted in Figures 24 and 25, $\lambda_{\max}^{b}$ continues to plateau near the instability threshold for two distinct step sizes, while $\lambda_{\max}$ once again settles at a lower level. Notably, as we vary the batch size, the gap between $\lambda_{\max}^{b}$ and $\lambda_{\max}$ increases for smaller batches, mirroring the patterns described in Section 5.

### v4

@@ Appendix I On Largest Eigenvalues of Sums of Matrices
I.1 Ordering the Largest Eigenvalues.
@@
I.2 Trends of $\lambda_{\max}$ given $b$
@@
As we discuss in Section J, while the behavior we observe for very small batches is not surprising, for bigger batch sizes it is. In computer vision tasks where there are way more parameters than datapoints, we observe that the gap between $\lambda_{\max}^{b}$ and $\lambda_{\max}$ decreases linearly with $1/b$ when evaluating the $\lambda_{\max}^{b}$ of any fixed model with different batch sizes, see Figure 22. This $1/b$ scaling is also what we expect from our mathematical analysis in Appendix I. We are indeed able to establish that for a fixed net, the gap scales as $1/b$ for small batch sizes, when the worst case $\lambda_{\max}^{b}$ is very different from $\lambda_{\max}$, and with $1/\sqrt{b}$ for big batch sizes [^13] Although big $b$ means such that the number of directions spanned in the parameter space by the vectors $\nabla_{\theta}f(\theta,x)$ are repeated multiple times, and that may be practically unrealistic with the current sizes of networks.. Indeed:
@@
where $\sigma^{2}=\frac{1}{b}\max\{\lVert\mathbb{E}[\sum_{i}M_{i}^{\top}M_{i}\rVert,\lVert\mathbb{E}[\sum_{i}M_{i}M_{i}^{\top}\rVert\}$ is the expected second moment of the matrices and $B\geq\lVert M_{i}-M\rVert$, for all $i$, is a bound to the biggest random matrix $M_{i}$.
I.3 Random Matrix Theory for Scaling Eigenvalues.
@@
Let $n_{1},n_{2},b\in\mathbb{N}$, let $M_{1},M_{2},\ldots,M_{b}\in\mathbb{R}^{n_{1}\times n_{2}}$ be independent random matrices satisfying $\mathbb{E}[M_{i}]=M$ and $\lVert M_{i}-M\rVert\leq B$ for all $i$, let $v=\max\{\lVert\mathbb{E}[\sum_{i}M_{i}^{\top}M_{i}\rVert,\lVert\mathbb{E}[\sum_{i}M_{i}M_{i}^{\top}\rVert\}$ then for all $t>0$
@@
Appendix J Dependence of $\lambda_{\max}^{b}$-$\lambda_{\max}$ Gap on the Batch Size
J.1 Highest Eigenvalue of Mini-Batch Hessian
For a batch size $b$, denote by
$\lambda_{\max}^{b}:=\mathbb{E}_{B\sim\mathcal{P}_{b}}\big{[}\lambda_{\max}(\mathcal{H}(L_{\mathbf{B}}))\big{]}.$
{mdframed}
[nobreak=true] Then, we establish here that:
  • Also $\lambda_{\max}^{b}$ stabilizes.
  • $\lambda_{\max}^{b}$ is always greater than $\lambda_{\max}$.
  • However, the gap between $\lambda_{\max}^{b}$ and $\lambda_{\max}$ typically goes as $1/b^{\alpha}$ generally with $\alpha=0.70$. This is surprising, as when fixing the network, the gap goes as $1/b$ or $1/b^{1/2}$. See Appendices I, J
Average $\lambda_{\max}^{b}$ for the mini-batch Hessians
We establish that $\lambda_{\max}^{b}$–the average over the possible batches sampled of the highest eigenvalue of the mini-batch Hessian–generally stabilizes at a value just bigger than $2/\eta$. This happens on a wide range or models and datasets. Refer to Appendix O
@@
Appendix K The Hessian and the Fisher Information Matrix Overlap
@@
Appendix L Exemplification Through a Simplified Models
A natural implicit notion of preceived curvature.
We give here a qualitative heuristic for which Batch Sharpness is a sensible notion of curvature along the step of SGD. When the optimization dynamics occur on a fixed loss landscape, as in GD or noise-injected GD, the stability criterion trivially reduces to examining a single Hessian. Stability in such cases hinges directly on properties of this single Hessian, typically characterized by $\lambda_{\max}$, or other norms. In contrast, mini-batch SGD inherently samples different landscapes–and consequently different Hessians–at every iteration. Thus, it is non-trivial to identify which statistical properties of these sampled Hessians govern the stability dynamics. Contrary to initial intuition, the average Hessian does not adequately characterize stability [^14] In Appendices L.1 and L.2 we provide two toy examples where is clear that stability has to depend on higher moments or different statistics of the mini-batch Hessians and gradients and not on the full-batch (averaged) quantities..
A natural first guess for what statistic reaches the value of $2/\eta$ would be a statistic of the maximum eigenvalue of the mini-batch Hessians, and not the maximum eigenvalue of the averaged Hessian. Indeed, if one has a few outlier Hessians, those may induce divergence even if the average Hessian is small. We also establish that is not the quantity to look at—and is generally bigger than Batch Sharpness—in Appendix O. While the landscape on the batch matter, the reason is that the (mis-)alignment of gradients and Hessians plays a bigger role. $\lambda_{\max}=2/\eta$ is the instability threshold if and only if the landscape is quadratic or the step is aligned with the eigenvector of the highest eigenvalue, here this is not the case. Batch Sharpness emerges naturally as the relevant measure because it explicitly captures the alignment between the gradient step direction and the curvature of the sampled mini-batch landscape at each iteration, effectively measuring the curvature perceived by SGD on each particular step.
L.1 Stability cannot Depend on Full-Batch Quantities—Quadratics
Let us now look at possible full-batch stability results, as developed by Cohen et al. [1]. If the right stability notion for mini-batch SGD depended on the full-batch Hessian or gradients, then it would be independent on $\gamma$, and this can not be the case. For instance, if a notion as the one in Equation (LABEL:eq:stab:2) would explain stability, then $\alpha$ could grow until $\alpha=2/\eta$. However, no matter the size of $\alpha$, if $\gamma>\max\big{\{}\alpha,\tfrac{2}{\eta}-\alpha\big{\}}$ a mini-batch iteration $x_{t+1}=(1-\eta\mathbf{A}_{i})x_{t}=$ increases the size of the second component of $x$ by a multiplicative factor bigger than 1.
Thus a study of the stability of the system can not depend on the full-batch Hessian but has to depend on how big the oscillations due to the size of $\gamma$ are, i.e., on the higher moments of the distribution of the mini-batch Hessian. Note that this situation would be even more extreme if we had (as in the practice of deep learning) the top eigenvectors of the mini-batch Hessians to point in completely different directions, not just to have high variance. As a sanity check, any stability threshold dependent on the higher moments of the mini-batch Hessian or gradients–as Batch Sharpness or $\mathbb{E}_{B}[\lambda_{\max}(\mathcal{H}(L_{\mathbf{B}})]$–would induce $\eta_{\max}$ to depend on $\gamma$ too.
L.2 Diagonal Linear Networks
@@
  • Smaller size: Thus setting $\lambda_{\max}$ equal to $\lambda$ means that the max between $\lambda_{1}$ and $\lambda_{2}$ is exactly $2\lambda$. Note that the fact that $a_{1}\cdot b_{1}=a_{2}\cdot b_{2}=1$ and Cauchy-Schwarz imply that $\lambda_{1},\lambda_{2}\geq 2$. Setting $\lambda_{\max}^{1}$ to $\lambda$ thus implies that the maximum between $\lambda_{1}$ and $\lambda_{2}$ is at most $22\lambda-$ 2, generally smaller.
@@
Appendix M Illustration of EoSS in Variety of Settings: Batch Sharpness
@@
We further confirm the EoSS regime in a five-layer CNN. As depicted in Figures 41 and 32, Batch Sharpness continues to plateau near the instability threshold for two distinct step sizes, while $\lambda_{\max}$ once again settles at a lower level. Notably, as we vary the batch size, the gap between Batch Sharpness and $\lambda_{\max}$ increases for smaller batches, mirroring the patterns described in Section 6.
@@
Figure 31: MLP: 2 hidden layers, hidden dimension 512; learning rate 0.01, 8k subset of CIFAR-10. Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue line).
@@
Figure 32: CNN: 5 layers (3 convolutional, 2 fully-connected), learning rate 0.03, 8k subset of CIFAR-10. Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue line).
@@
Figure 33: ResNet-10, learning rate 0.005, 8k subset of CIFAR-10. Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue line).
Appendix N Illustration of EoSS for the SVHN dataset
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 34: MLP: 2 hidden layers, hidden dimension 512; learning rate 0.01, 8k subset of SVHN. Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue line).
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 35: CNN: 5 layers (3 convolutional, 2 fully-connected), learning rate 0.05, 8k subset of SVHN. Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue line).
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 36: ResNet-10, learning rate 0.005, 8k subset of SVHN. Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue line).
Appendix O Illustration of EoSS in Variety of Settings: $\lambda_{\max}^{b}$
@@
To assess whether increased depth alters the phenomenon, we use a deeper MLP (MLP_L) with four hidden layers, training again on the same CIFAR-10 subset. Figures 39 and 40 show that $\lambda_{\max}^{b}$ exhibits the same EoSS behavior for two different step sizes, reinforcing that depth alone does not invalidate our findings.
@@
We further confirm the EoSS regime in a five-layer CNN. As depicted in Figures 41 and 42, $\lambda_{\max}^{b}$ continues to plateau near the instability threshold for two distinct step sizes, while $\lambda_{\max}$ once again settles at a lower level. Notably, as we vary the batch size, the gap between $\lambda_{\max}^{b}$ and $\lambda_{\max}$ increases for smaller batches, mirroring the patterns described in Section 6.